In a class of 125 students 20% of the students can dance.2/5th of the total number of students in the class can sing and 2/5 of the remaining number of students in the class are good at sports.What is the respective ratio of the number of students who can dance to the number of students who are good at sports?
Found 2 solutions by geetha_rama, lwsshak3:
Answer by geetha_rama(94)

No: of students who can dance = (20/100)*125 = 25
No: of students who sings = (2/5)*125 = 50
No: of students good at sports = (2/5) * (125-25-50) =20
Ratio of students who can dance to that of sports = 25:20 = 5:4

Answer by lwsshak3(11628)

In a class of 125 students 20% of the students can dance.2/5th of the total number of students in the class can sing and 2/5 of the remaining number of students in the class are good at sports.What is the respective ratio of the number of students who can dance to the number of students who are good at sports?
***
20%*125=25=students can dance
2/5*125=50=students can dance
ratio of the number of students who can dance to the number of students who are good at sports? 25/50=1/2